# A Calculus of Overlays

### v1

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, Overlay-Calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics builds on only naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance simply does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence. Overlay-Calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are CPS-agnostic, records natively encode random-access memory, and self-reference resolves to multiple targets. These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves remarkable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself—Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so powerful.
@@
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures—in the spirit of F-coalgebras (Rutten, 2000)—where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006)—over 100,000 packages—behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation and by the NixOS module system’s recursive merging mechanism, we built the Overlay language (Yang, 2025b)—a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
@@
In developing real programs in the Overlay language—Church-encoded data types (Appendix B), an Expression Problem solution (Section 5), trie-based random-access memory (Appendix A), and continuation patterns (Appendix G)—we discovered that several long-standing problems in programming language theory do not arise. The linearization problem that has plagued mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) does not exist—not because it is resolved, but because the three primitives make it structurally impossible (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. These are not properties we engineered; they are phenomena we observed in working programs.
This paper distills these observations into Overlay-Calculus, a minimal computational model for declarative programming. It turns out that function application has a straightforward translation into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into Overlay-Calculus, Overlay-Calculus is Turing complete. The translation reveals an asymmetry: while embedding the $\lambda$-calculus into Overlay-Calculus is trivial, the reverse is non-trivial because Overlay-Calculus provides primitive operations—self-referential records, open recursion, symmetric inheritance—that have no direct counterparts in the $\lambda$-calculus. Moreover, Overlay-Calculus naturally encodes random-access memory: records are tries, and inheritance is trie union, yielding an immutable analogue of the RAM machine (Appendix A). This asymmetry suggests that Overlay-Calculus is not merely a syntactic variant of the $\lambda$-calculus but captures different computational patterns.
@@
All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example. All tests are mechanically verified against expected output snapshots.
@@
  • Inheritance operates at the tree level: a record is a set of elements—definitions and inheritance sources—so inheriting from two overlays that share a label recursively merges their subtrees (union mount semantics), rather than overriding at the method level as in mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi simply does not arise.
@@
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree—a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala, Appendix C; NixOS module system, Appendix D). Since inheritance is idempotent, all objects contribute equally—no disambiguation is needed.
  • The $\lambda$-calculus (in A-normal form) translates into Overlay-Calculus in five rules, so Overlay-Calculus is Turing complete. Composing this translation with the semantic functions induces a semantics on the $\lambda$-calculus that requires no function spaces, no substitution, and no reduction rules—only set membership in powerset lattices (Section 4). The translation is adequate and fully abstract with respect to Böhm tree equivalence (Section 4.2): two terms have the same overlay-convergence behavior in all contexts if and only if they have the same Böhm tree. This yields a fully abstract semantics of the lazy $\lambda$-calculus that lives entirely in naive set theory—powerset lattices and the Knaster–Tarski fixed-point theorem—without domain theory, inverse limits, or Scott continuity. Standard data types (booleans, natural numbers) can be Church-encoded without scalars (Appendix B).
  • We observe several emergent properties in practice: the Expression Problem (Wadler, 1998) does not arise (Section 5), programs are CPS-agnostic (Appendix G), and records natively encode random-access memory (Appendix A).
Section 5 demonstrates these properties through a case study of the Expression Problem.
@@ 2. Syntax
An expression is a set of elements enclosed in braces. Each element is either a definition $\ell\mapsto e$ or an inheritance source. Because expressions are sets, element order is irrelevant and duplicates have no effect—inheritance is inherently commutative, idempotent, and associative.
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in Overlay-Calculus. The Overlay language also provides scalar types through a foreign-function interface (FFI); Overlay-Calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI—scalars simply did not arise.
@@
  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n$ is a de Bruijn index (de Bruijn, 1972) counting the number of enclosing scope levels to traverse.
Both forms retrieve the target scope—the fully inherited record after all inheritance has been applied—and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it. The named form desugars to the indexed form during parsing (Section 3).
A new scope level is created at each $\{\ldots\}$ (record literal). The scope contains all properties of that overlay, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single overlay whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ scope levels above) retrieves the enclosing record named $\mathrm{Foo}$.
@@ 3. Overlay Trees
[…]
@@
  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index (de Bruijn, 1972) (zero-based) and $\ell_{\mathrm{down},*}$ is the list of downward projections.
During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$. The indexed form $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ already carries the de Bruijn index $n$ directly; no resolution is needed. A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$. Let $p_{\mathrm{def}}$ be the prefix of $p$ up to and including that occurrence. The de Bruijn index is $n=|p|-|p_{\mathrm{def}}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$. The de Bruijn index is $n=|p|-|p^{\prime}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
@@
The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overlays}$ first.
@@
Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$. Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
@@
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references—analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
@@
Intuitively, $\mathtt{resolve}$ turns a syntactic reference into the set of paths it points to in the fully inherited tree. It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$. Resolution proceeds in two phases: first, the qualified this resolution function $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
@@
When multiple routes exist (due to multi-path inheritance), they may yield different target paths; all are collected. We now define $\mathtt{this}$.
@@
Intuitively, $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” The design of $\mathtt{this}$ reconciles two historically opposing approaches to scope resolution. Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack; this is flexible but unpredictable. Scheme (Sussman and Steele Jr., 1975) corrected this with lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid. Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path. When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix C) and the NixOS module system rejects the duplicate option declaration (Appendix D)—neither can express that all routes are equally valid.
Overlay-Calculus combines the strengths of both approaches. The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable. But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one. All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless. The only observable is property presence, making multi-path resolution well-defined without disambiguation.
Concretely, each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\small\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
@@
Alternative approaches considered
Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple inheritance routes to the same enclosing scope.
In retrospect, the difficulty was inevitable. A closure captures a single environment, so a fixed-point combinator solves for a single self-reference; when multiple routes exist, it attempts to unify them into one value. A stack-based environment is a linear chain—each scope level has one parent—so looking up a de Bruijn index returns a single value. Both data structures embed a single-valued assumption that is incompatible with the multi-path situation that arises when independent overlays inherit from the same enclosing scope (Appendix C). Equation (6) emerged from recognizing this: it operates on paths as first-order data and tracks a frontier set $S$ rather than a single current scope.
@@
The definitions above constitute the complete semantics of Overlay-Calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. This is a deliberate design choice, not an omission. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In Overlay-Calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything—it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument—identifying the lattice, verifying positivity of each equation, and handling the transitive closure $\mathtt{bases}^{*}$—is given in Appendix E. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point—analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.
4. Translation from $\lambda$-Calculus

### v2

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984). Overlay-calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are function color blind (Nystrom, 2015), ordinary arithmetic yields the relational semantics of logic programming, and self-reference resolves to multiple targets, making overlay-calculus strictly more expressive than the $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991). These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
@@
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, in the spirit of F-coalgebras (Rutten, 2000), where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. Through iterative refinement, beginning with an early version implemented in Nix, the system evolved into the Overlay language (Yang, 2025b), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
@@
In developing real programs in the Overlay language, including boolean logic and natural number arithmetic (Section 5) and a web application (Section 5), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. We also found that boolean logic and arithmetic, implemented as separate overlay files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (Section 5); and that the same web application overlay file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section 5).
This paper distills these observations into overlay-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into overlay-calculus, overlay-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into overlay-calculus requires five local, syntax-directed rules (Section 4), but the reverse is impossible—overlay enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
@@
All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example. [^1] The Overlay language is statically typed and therefore stricter than the untyped overlay-calculus. All examples in this paper are ported from the Overlay language and conform to both semantics. Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
@@
  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two overlays that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi does not arise.
@@
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B). Since inheritance is idempotent, all objects contribute equally; no disambiguation is needed.
  • The $\lambda$-calculus (in A-normal form) translates into overlay-calculus in five rules (Section 4). The translation is fully abstract with respect to Böhm tree equivalence (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section 5).
  • Open extension via overlay changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express overlay (Section 4.2).
Section 5 demonstrates these properties through a case study of Nat arithmetic.
@@ 2. Syntax
An expression is a set of elements enclosed in braces. Each element is either a definition $\ell\mapsto e$ or an inheritance source. Because expressions are sets, element order is irrelevant and duplicates have no effect; inheritance is inherently commutative, idempotent, and associative.
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in overlay-calculus. The Overlay language also provides scalar types through a foreign-function interface (FFI); overlay-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
@@
  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
Both forms retrieve the target scope, that is, the fully inherited record after all inheritance has been applied, and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it. The named form desugars to the indexed form during parsing (Section 3).
A new scope level is created at each $\{\ldots\}$ record literal. The scope contains all properties of that overlay, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single overlay whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ steps above the enclosing scope) retrieves the record named $\mathrm{Foo}$.
@@ 3. Overlay Trees
Unlike the $\lambda$-calculus, which requires reduction rules ($\beta$-reduction) to specify how terms compute, overlay-calculus has no reduction steps. There is no rewriting of terms, no evaluation order, and no notion of a “result” that a term reduces to. Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree. The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics. This is the observational perspective introduced in Section 1: like an F-coalgebra (Rutten, 2000), an overlay tree’s behavior is determined by observation (property presence), not by reduction.
@@
  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index and $\ell_{\mathrm{down},*}$ is the list of downward projections.
During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$. The indexed form $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ already carries the de Bruijn index $n$ directly; no resolution is needed. A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$. Let $p_{\mathrm{target}}$ be the prefix of $p$ up to and including that occurrence. The de Bruijn index is $n=|p|-|p_{\mathrm{target}}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$. The de Bruijn index is $n=|p|-|p^{\prime}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
@@
Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$. The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overlays}$ first.
@@
Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
@@
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
@@
Intuitively, $\mathtt{resolve}$ turns a syntactic reference into the set of paths it points to in the fully inherited tree. It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$. Resolution proceeds in two phases: $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
@@
When multiple routes exist due to multi-path inheritance, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
@@
$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
@@
The definitions above constitute the complete semantics of overlay-calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In overlay-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.
4. Embedding the $\lambda$-Calculus